A chemist titrates 130.0 mL of a 0.0647 M methylamine (CH3NH2 solution with 0.5334 M HNO3 solution at 25 °C. Calculate the pH at equivalence. The p K, of methylamine is 3.36 Round your answer to 2 decimal places Note for advanced students: you may assume the total volume of the solution equals the initial volume plus the volume of HNO3 solution added pH =
Answer #1

pKb = 3-36 mmoles of CH3NH2-130xo-0647 No ef mmoles ef BD577x 33 130ΧΟ,064าะ ½XO, 5334= 15-11m) = 8.HI 14 5 구구 - 7-1-06